I need to solve this by addition method and determine whether they are independent, dependent, or inconsistent.
Here is the equation:
-3x+2y=8
3x+2y=8

Answer by elima(1433)

-3x+2y=8
3x+2y=8
---------
4y=16
y=4
==========
-3x+2(4)=8
-3x+8=8
-3x=0
Undefined, so it will be inconsistent
:)